### v1

# SEEA-R1: Tree-Structured Reinforcement Fine-Tuning for Self-Evolving Embodied Agents

Abstract
00footnotetext: * Equal contribution.00footnotetext: $\dagger$ Project leader: jason.ju@x-humanoid.com00footnotetext: $\ddagger$ Corresponding authors: shanghang@pku.edu.cn, jian.tang@x-humanoid.com
[…]
Figure 1: SEEA-R1 self-evolves by planning with perception and tree-based search, refining actions via reward signals to solve tasks from high-level instructions in embodied environments.
1 Introduction
Embodied agents commonly operate in complex, long-horizon environments that require not only low-level perception and motor control, but also high-level reasoning, planning, and decision-making capabilities. Despite recent advances in large language models (LLMs) guo2025deepseek ; shao2024deepseekmath ; jin2025search and multi-modal LLMs (MLLMs) liu2025visual ; zhan2025vision ; huang2025vision ; yu2025perception have greatly improved agents’ abstract reasoning and perception, extending these abilities to open-ended embodied settings remains a fundamental challenge. Since embodied agents often operate interactively over long-horizon tasks, constantly perceiving visual feedback to produce actions, existing LLMs and MLLMs fall short: LLMs often lack perceptual grounding zhang2025embodiedreasonersynergizingvisualsearch , while MLLMs continue to struggle with structured multi-step planning, maintaining long-term coherence, and adapting to dynamic environments huang2025vision ; salim2025viper .
To tackle general-purpose embodied intelligence, we argue that agents must acquire self-evolution—the capacity to self-generate training signals and refine reasoning through closed-loop learning. By continuously interacting with the environment and learning from their own experiences, self-evolving embodied agents can bridge the gap between perception and cognition, generalize to novel tasks, and develop long-horizon planning capabilities that transcend the limitations of static, supervised learning paradigms.
In this context, reinforcement fine-tuning (RFT) guo2025deepseek ; shao2024deepseekmath has emerged as a promising paradigm: By providing reward-based feedback, RFT enables agents to learn from trial-and-error trajectories and gradually refine their decision-making policies. However, bridging RFT to embodied agents introduces unique challenges that are absent in symbolic domains such as mathematics or code generation. In this work, we identify two major challenges for reinforce fine-tuning agents in embodied environments: 1) Lack of intermediate feedback in multi-step reasoning tasks: In embodied tasks with long horizons and delayed sparse rewards, it is difficult to assign credit to intermediate decisions, making it hard for RFT to effectively guide policy learning. 2) Poor generalization of handcrafted reward functions: Existing RFT pipelines typically rely on simulator-specific or task-specific reward signals that do not generalize to novel environments, making it hard for agents to self-improve across diverse tasks.
To address these challenges, we introduce SEEA-R1 (Self-Evolving Embodied Agents), the first framework to adapt RFT for training embodied agents capable of long-horizon reasoning and autonomous self-evolution. SEEA-R1 integrates two key components. 1) Tree-GRPO (Tree-based Group Relative Policy Optimization): we extend Group Relative Policy Optimization with Monte Carlo Tree Search (MCTS), enabling agents to explore alternative trajectories and convert sparse outcome rewards into structured, step-wise process signals. This improves credit assignment and facilitates reasoning over extended action sequences. 2) MGRM (Multi-modal Generative Reward Model): To eliminate the reliance on handcrafted or environment-specific reward functions, we introduce MGRM—a reward model trained from multi-modal, multi-turn trajectories. MGRM provides task-agnostic reward estimation, supporting generalization and self-improvement across diverse embodied scenarios.
We evaluate SEEA-R1 on the ALFWorld benchmark, which rigorously tests an agent’s planning and reasoning capabilities by requiring it to map abstract goals to visually grounded action sequences. Our proposed method achieves state-of-the-art success rates of 85.07% and 36.19% on textual and multi-modal tasks, respectively, outperforming previous models including Qwen2.5-VL and GPT-4o. To further evaluate self-evolution in a realistic yet challenging setting, we replace ground-truth rewards with MGRM-based self-supervised signals; SEEA-R1 still achieves 80.30% and 23.88% on textual and multi-modal tasks, respectively, surpassing all open-source baselines. These results highlight the effectiveness and scalability of SEEA-R1 as a self-evolving embodied agent.
Our main contributions are as follows:
(1) We propose SEEA-R1, the first reinforcement fine-tuning framework designed to support self-evolving capabilities in embodied agents. (2) By introducing Tree-GRPO, we augment GRPO with MCTS to enable dense and interpretable credit assignment across multi-step trajectories, (3) We replace handcrafted reward signals with MGRM, a multi-modal generative reward model that rewards task completion. (4) We achieve new state-of-the-art performance on the ALFWorld benchmark under both supervised and self-supervised settings, demonstrating strong planning, reasoning, adaptability, and generalization. (5) To facilitate future research and applications in the embodied intelligence community, we will open-source our full framework and modular components—including our reward model MGRM, and training pipelines.
2 Related Works
2.1 Self-Evolution Learning
Recent research has explored how large language and vision-language models can self-evolve through mechanisms such as self-play, reinforcement learning, and trajectory refinementxu2024survey . Welcome to the Era of Experience silver2025welcome highlights the importance of learning from experience and real-world interactions, emphasizing two key components: reward mechanism for environmental interaction and long-horizon perception-action task loop. SPC chen2025spcevolvingselfplaycritic utilizes an adversarial self-play environment to enhance reasoning, while rStar-Math guan2025rstar and Agent-Q putta2024agentqadvancedreasoning demonstrates that small models can improve using MCTS with reward modeling. Dreamsmooth lee2023dreamsmooth and R2I samsami2024mastering incorporate planning and prediction modeling into reward design to address long-horizon tasks. RAGEN wang2025ragen introduces a modular system to enhance LLM reasoning capabilities in multi-turn interactive environments. Building on these foundations, our approach integrates MCTS into the self-evolution loop, enabling action rollout, trajectory filtering, and policy refinement. We scale it for the first time to vision-language decision-making in long-horizon, physically grounded tasks.
2.2 Planning for Embodied Agent
LLM-Planner song2023llm , LLM+P liu2023llm+ , and RegionFocus luo2025visualtesttimescalinggui have demonstrated that large language and vision-language models can significantly enhance task performance. Planning is also critical for embodied agents to make sequential decisions and achieve long-term goals. Training-free approaches such as PIVOT nasiriany2024pivotiterativevisualprompting and ECoT zawalski2024robotic leverage the spatial reasoning capabilities of VLMs to plan effectively, leading to improved task success rates in embodied agents. In contrast, training-based methods like VIPER salim2025viper and MPO xiong2025mpo combine vision-language perception with LLM-based reasoning or meta-planning, but often depend on predefined modules or expert-designed strategies. Other works such as ARMAP chen2025scaling and LS-Imagine li2024open show that incorporating the planning module into the reward assignment process during training can significantly improve long-horizon task performance in web search and Minecraft environments. In the embodied domain, EmbodiedEval cheng2025embodiedeval highlights persistent planning failures in current embodied systems, indicating a gap between perception, reasoning, and decision execution. To address this, our approach integrates MCTS-based planning directly into training in embodied scenarios, which, to our knowledge, has not been previously explored in this context.
2.3 Reinforcement Fine-Tuning
DeepSeek-R1 guo2025deepseek  muennighoff2025s1simpletesttimescaling shows that RL with formatting and result-only rewards can steer LLMs toward humanlike, complex chain-of-thought reasoning, boosting performance on challenging tasks. Subsequently, Search-R1 jin2025search , TORL li2025torl explored extending the LLM-R1 approach to tool use. R1 approach has also achieved success in the visual domain. Vision-R1 huang2025vision , Visual-RFT liu2025visual , and Video-R1 zhan2025vision improve visual grounding through reward-based reinforcement fine-tuning, while R1-OneVision yang2025r1 and Perception-R1 yu2025perception formalize visual input into language for structured reasoning. In the embodied AI domain, robotxR1 boyle2025robotxr1 transforms visual information into textual descriptions to train LLMs, enhancing their robotic scenario decision-making capabilities. Embodied-R zhao2025embodied further extends to spatial reasoning. But they are not all trained as multimodal models and still rely on task-specific priors. To address these limitations, we developed the first general embodied R1 framework SEEA-R1, which integrates a vision-grounded reward model with MCTS-based planning to enable self-supervised policy optimization without handcrafted rewards.
Figure 2: SEEA-R1 framework. The framework drives continuous improvement through an iterative loop of two core cycles as follows: 1.Data Evolution: The Policy Model interacts with the environment via MCTS from an initial state to generate the experience dataset, containing trajectories with derived Q-values, ground truth rewards from the environment, and rewards from the current Reward Model. 2.Model Evolution: The collected data is used to update both models: (a) The Policy Model to predict actions and (b) The Reward Model to predict categorical outcomes. Refined models from Model Evolution then drive the next Data Evolution iteration, enabling continuous self-evolution.
3 Methods
In this section, we first formulate the Embodied Agent, the Self-Evolving Embodied Agent, and Multi-round Interaction Embodied Scenario. Then, we briefly introduce the MCTS coulom2006efficient algorithm. Next, we present SEEA-R1, detailing how data evolution and model evolution are implemented. Finally, we summarize the overall self-evolution loop.
3.1 Preliminaries
Embodied Agent
in our methods is defined as a 4-tuple $(s_{t},a_{t},r_{t},o_{t})$ at timestep $t$. It interacts with the environment by receiving observations $o_{t}$ and taking actions $a_{t}$ in each step, with state $s_{t}$ maintaining the full interaction history, and reward $r_{t}$ reflecting task progress. Further details can be found in Appendix C. A Self-Evolving Embodied Agent is defined to satisfy three core conditions: automated data synthesis, iterative capability improvement, and closed-loop reasoning system he2025survey . Multi-round Interaction Embodied Scenario is modeled as a Partially Observable Markov Decision Process (POMDP) drake1962observation ; aastrom1965optimal , formalized as a 7-tuple $(\mathcal{S},\mathcal{A},T,R,\mathcal{O},\Omega,\gamma)$. Here, $\mathcal{S}$ denotes the partially observable state space, $s_{t}\in\mathcal{S}$; $\mathcal{A}$ is the action space, $a_{t}\in\mathcal{A}$; $T(s_{t+1}|s_{t},a_{t})$ represents the transition probability; $r_{t}$ is the reward function, $r_{t}\in R$; $\mathcal{O}$ is the observation space, $o_{t}\in\mathcal{O}$; $\Omega(o_{t+1}|s_{t},a_{t})$ is the observation probability; and $\gamma$ is the discount factor.
Monte Carlo Tree Search (MCTS) coulom2006efficient
is a heuristic algorithm for decision-making in large state spaces widely used in planning, games, and embodied reasoning tasks. After initialization, MCTS iteratively performs four steps: Selection: Starting from the root, the agent selects child nodes with the highest UCT scores until reaching a leaf node; Expansion: The agent executes the selected action $a_{L}$ to obtain observation $o_{L}$, constructs a new node $s_{L+1}=\{s_{L},(a_{L},o_{L})\}$, and expands it with a set of candidate natural language actions $\{a_{L+1,i}\}_{i=1}^{G}$; Simulation: From $s_{L+1}$, multiple rollouts are simulated until termination, giving up, or reaching the depth limit, yielding trajectories $s_{0}\rightarrow\dots\rightarrow s_{T}^{(j)}$; Backup: Rollout results are used to update visit counts $N(s_{t},a_{t})$, cumulative rewards $R^{(j)}(s_{t},a_{t})$, and action values $Q(s_{t},a_{t})$ along each path back to the root. Further details are provided in Appendix D.
3.2 Self-Evolving Emodied Agent-R1
As illustrated in Figure 2, SEEA-R1 continuously improves the agent through two iterative cycles: Data Evolution, where the agent to interact with the environment via MCTS to generate experiential data, and Model Evolution, where both the Policy Model and Reward Model are updated using this data. The evolved models are then used for the subsequent Data Evolution, forming a closed self-improvement loop. The details of each phase are provided below.
3.2.1 Data Evolution: Experience Generation via MCTS
Figure 3: Monte Carlo Tree Search (MCTS) in SEEA-R1. (a) Selection Traverse tree via UCT (2) until reaching a leaf. (b) Expansion Execute action, observe result, and expand with new actions. (c) Simulation Roll out from new node to termination or depth limit, collecting reward $r$. (d) Backup Propagate rewards to update action values $Q$ using the formulation in Equation (3).
To convert sparse outcome rewards into dense process rewards for multi-step reasoning, SEEA-R1 employs Monte Carlo Tree Search (MCTS) for experience generation. MCTS estimates action-values, $Q(s_{t},a_{t,i})$—the expected future reward for action $a_{t,i}$ in state $s_{t}$—which serve as process rewards. This provides immediate, step-by-step feedback from otherwise sparse environmental signals. The precision of these Q-values and therefore the quality of the rewards of the process improves as MCTS performs more simulations (Figure 3) exploring a wider range of trajectories. This progressive refinement of MCTS-generated experiential data is how SEEA-R1 implements ’Data Evolution’, leading to increasingly reliable learning guidance for the agent.
3.2.2 Model Evolution: Co-refining Policy Model and Reward Model
Multi-modal Generative Reward Model (MGRM) Training.
[…]
Embodied Agent Policy Update through Tree-GRPO.
Considering the Embodied Agent as a parameterized model $\pi_{\theta}(a_{t}|s_{t})$ with parameters $\theta$, which generates a set of available actions $\{a_{t,i}\}_{i=1}^{G}$ based on the current state $s_{t}$, the GRPO shao2024deepseekmath loss can be reused for this group of actions at each node of the tree-structured experiential data:
$\mathcal{J}(\theta)=\,$ $\mathbb{E}_{\begin{subarray}{c}s_{t}\sim\mathcal{S},\{a_{t,i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\mathcal{A}|s_{t})\end{subarray}}$
$\Biggl{[}\frac{1}{\sum_{i=1}^{G}\lvert a_{t,i}\rvert}\sum_{i=1}^{G}\sum_{k=1}^{\lvert a_{t,i}\rvert}\min\bigl{(}\rho_{t,i,k}(\theta)\hat{A}_{t,i,k},\mathrm{clip}\bigl{(}\rho_{t,i,k}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}}\bigr{)}\hat{A}_{t,i,k}\bigr{)}\Biggr{]}$
$-\beta\mathbb{D}_{\mathrm{KL}}\bigl{[}\pi_{\theta}(\cdot|s_{t})\|\pi_{\mathrm{ref}}(\cdot|s_{t})\bigr{]}$ (1)
where $\rho_{t,i,k}(\theta)=\frac{\pi_{\theta}(a_{t,i,k}|s_{t},a_{t,i,<k})}{\pi_{\theta_{\text{old}}}(a_{t,i,k}|s_{t},a_{t,i,<k})},\,\hat{A}_{t,i,k}=\frac{pr_{t,i}-mean(\{pr_{t,i}\}_{i=1}^{G})}{std(\{pr_{t,i}\}_{i=1}^{G})}$
Here, $a_{t,i,k}$ denotes the $k$-th token of the $i$-th action $a_{t,i}$ at timestep $t$, and $a_{t,i,<k}$ denotes the tokens preceding token $k$ in action $a_{t,i}$. $\pi_{\theta}$ is the current Embodied Agent, $\pi_{\theta_{old}}$ is the old Embodied Agent that collected the experiential data, and $\pi_{\mathrm{ref}}$ is the initial Embodied Agent. $\rho_{t,i,k}(\theta)$ represents the importance sampling ratio. $\hat{A}_{t,i,k}$ is the advantage function for the $k$-th token of the $i$-th action at timestep $t$. $\epsilon_{\mathrm{high}}$ and $\epsilon_{\mathrm{low}}$ are from the Clip-Higher strategy of DAPO yu2025dapo . $\mathbb{D}_{\mathrm{KL}}\bigl{[}\pi_{\theta}(\cdot|s_{t})\|\pi_{\mathrm{ref}}(\cdot|s_{t})\bigr{]}$ is the k3 estimator of KL divergence between $\pi_{\theta}$ and $\pi_{\mathrm{ref}}$. The coefficient $\beta$ controls its weight.
3.2.3 Iterative Self-Evolution Loop
The SEEA-R1 framework iteratively alternates between Data Evolution and Model Evolution. In each cycle, the updated Embodied Agent and MGRM from the Model Evolution phase are deployed in the subsequent Data Evolution phase. This iterative evolution process, detailed in Appendix, allows the SEEA-R1 agent to progressively enhance its reasoning, reward understanding, and task execution capabilities over time.
4 Experiments
This section presents an experimental evaluation of the effectiveness of our proposed framework on embodied tasks. The section is structured as follows. We begin by outlining the experimental setup and implementation details. Then, we demonstrate our experimental results comparisons with existing baselines, followed by an in-depth analysis and ablation studies to examine the effects of different training algorithms and our self-evolution reward design.
4.1 Experimental Setup
4.1.1 Environment, Dataset and Evaluation
ALFWorld. We train and evaluate our approach on ALFWorld shridhar2020alfworld , an interactive simulation environment and a widely used benchmark for evaluating embodied agents on complex, long-horizon tasks in simulated household environments. We leverage this interactive benchmark to collect data for model training, using the rewards provided by ALFWorld as our ground truth (GT) rewards. ALFWorld provides two types of observations: textual descriptions and raw visual inputs (images), enabling evaluation in both language-only and vision-language settings. Further details on the dataset and expert trajectory collection can be found in Appendix.
EmbodiedEval. To evaluate the generalization ability of our embodied agents beyond the training environment, we introduce EmbodiedEval cheng2025embodiedeval as an out-of-distribution benchmark. EmbodiedEval tests MLLMs as embodied agents across diverse tasks—including Attribute Question Answering (AttrQA), Spatial Question Answering (SpatialQA), Navigation, Object Interaction, and Social Interaction—within 125 realistic 3D scenes. It provides a comprehensive assessment of agent capabilities in previously unseen scenarios. This setup enables us to measure generalization under significant domain shifts compared to the ALFWorld environment. We analyze the impact of different control strategies on performance, using key metrics such as overall accuracy, which reflects the percentage of fully completed tasks.
Evaluation.
To ensure experimental reproducibility, we set the decoding temperature to zero. When generating task completion trajectories, we include a one-shot in-context example for each task. Detailed prompts are provided in Appendix. Our primary evaluation metric is the Average Success Rate, which calculates the mean Success Rate across all test set task instances.
4.1.2 Implementation Details
MCTS
SEEA-R1 is employed to sample experience trajectories with process rewards for training. Each MCTS performs 30 search iterations, with each simulation proceeding up to a maximum depth of 30 steps. During the expansion phase, up to five child nodes—corresponding to candidate actions—are expanded from each selected leaf node.
Training Settings
We use the Qwen2.5-VL-7B-Instruct qwen2.5-VL as the base model to build our embodied agents. Since experience learning in our framework consists of two stages—trajectory sampling and model training—we define each iteration as one complete cycle of these two steps. To ensure fair comparison across different training paradigms (SFT, DPO and GRPO), we use consistent sample data sizes and training hyperparameters. For each iteration, the model is trained on newly collected data for one epoch as steps is 4. The total number of iterations is not fixed and proceeds until convergence. Specifically, for Tree-GRPO, model updates are performed after collecting 512 valid samples (i.e., with non-zero advantage values). For SFT, 512 trajectories with positive advantages are used. For DPO, 512 positive-negative trajectories pairs are sampled from the same parent node in the MCST. When training the multimodal generative reward model (MGRM) via GRPO, the group size is set to 24. All training adopt the same hyperparameters: a cosine annealing learning rate schedule (initial LR: 1e-6, warmup ratio: 0.05), batch size of 128, and KL divergence coefficient $\beta$ of 0.01. Experiments are conducted on 8 NVIDIA A100 80GB GPUs using the ms-swift framework zhao2024swiftascalablelightweightinfrastructure .
4.2 Experimental Results
4.2.1 Overall Performance Comparison
Table 1: Comparison of MLLM methods on unseen tasks (*-reported in previous work. This convention is also used for other tables.). Success rate (0-100 scale) is shown per task and overall.
Model Parameter Size Optimization Policy ALFWorld Unseen Avg
Pick Clean Heat Cool Look Pick2
GPT-4o achiam2023gpt - - 44 22 29 27 7 23 24
Florence-2* xiao2024florence 0.77B SFT 0 0 0 0 6 0 1
Idefics-2* laurenccon2024matters 8B SFT 4 0 0 0 6 0 2
MiniGPT-4* zhu2023minigpt 7B SFT 4 0 19 17 17 16 16
InstructBLIP* instructblip 7B SFT 50 26 23 6 6 0 22
Qwen2.5-VL qwen2.5-VL 7B SFT + DPO 6 21 11 22 5 4 11
RL4VLM* zhai2024fine 7B PPO 47 10 14 18 14 18 21
SEEA-R1 7B Tree-GRPO 49 40 43 41 24 16 36
Our proposed SEEA-R1 method, utilizing Qwen2.5-VL-7B-Instruct and Qwen2.5-7B-Instruct as respective base models, was evaluated against strong baselines on ALFWorld test set. As reported in Tables 1 and 2, SEEA-R1 was tested in two settings: text+visual (MLLM), with image observations, and text-only (LLM), with textual descriptions. In the more challenging MLLM setting requiring direct visual perception, SEEA-R1 achieved a 36.19% success rate, substantially outperforming GPT-4o and other leading open-source models, and demonstrated superior efficiency (23.37 vs. 25.10 average steps for GPT-4o). Furthermore, in the text-only (LLM) setting, SEEA-R1 attained an 84.3% success rate, with average completion steps (14.75) on par with GPT-4o.
Table 2: Performance comparison on the ALFWorld dataset. ITO stands for Inference-Time Optimization. MPO (Meta Plan Optimization) xiong2025mpo provides the agent with the meta plan as context in inference time to improve performance.
Model w/o Training ITO ALFWorld Average
Seen Unseen
GPT-4o* achiam2023gpt - 78.6 83.6 81.1
Qwen2.5-7B-Instruct* Yang2024Qwen25TR 71.4 75.4 73.4
Llama-3.1-8B-Instruct* grattafiori2024llama - 22.9 28.4 25.7
Llama-3.1-70B-Instruct* grattafiori2024llama - 78.6 73.9 76.3
GPT-4o* MPO 89.3 93.3 91.3
Llama-3.1-8B-Instruct* MPO 50.0 52.2 51.1
Llama-3.1-8B-Instruct + SFT* zeng2023agenttuning - 79.3 71.6 75.5
Llama-3.1-8B-Instruct + ETO* song2024trial - 77.1 76.4 76.8
Llama-3.1-8B-Instruct + KnowAgent* zhu2024knowagent - 80.0 74.9 77.5
Llama-3.1-8B-Instruct + WKM* qiao2024agent - 77.1 78.2 77.7
Llama-3.1-8B-Instruct + SFT* MPO 80.7 81.3 81.0
Llama-3.1-8B-Instruct + ETO* MPO 85.0 79.1 82.1
SEEA-R1 - 85.3 85.1 85.2
4.2.2 Comparative Study of Training Algorithms
To validate our proposed Tree-GRPO, we compared it against MCTS combined with DPO and SFT on ALFWorld. As shown in Figure 4, Tree-GRPO (blue line) consistently achieves superior task success rates (Figure 4a) and greater efficiency with fewer average steps (Figure 4b) than MCTS + DPO (purple line) and MCTS + SFT (red line). This demonstrates the enhanced performance of integrating GRPO within our MCTS-driven self-evolving framework for long-horizon embodied tasks.
4.2.3 Ablation Study: Self-Evolution with MGRM
We conducted an ablation study to demonstrate that our SEEA-R1 can achieve sustained performance improvement by leveraging its internally trained Multi-Modal Generative Reward Model (MGRM), thereby reducing reliance on ground truth (GT)rewards from the simulator. We compared three critical configurations with results shown in Figure 5:
Figure 4: Performance comparison of SEEA-R1 using different optimization algorithms on ALFWorld over training iterations, more detailed figures are provided in Appendix.
  1. 1.
    SEEA-R1 with GT Rewards: Policy optimization is guided directly by GT environment rewards. As expected, this (typically the blue line) yields the highest success rate, establishing an upper bound due to the perfectly accurate reward signal.
  2. 2.
    SEEA-R1 with frozen MGRM: An untrained and frozen MGRM provides reward signals for policy optimization. Notably, the orange line shows a high initial success rate in the first iteration despite the inaccurate reward signal. It’s possible that the initial exploration coincidentally aligned with the flawed reward. Nevertheless, the fundamental unreliability of the frozen MGRM ultimately misguides the learning process, leading to performance degradation in later iterations.
  3. 3.
    SEEA-R1 with training MGRM: Our proposed method, where the MGRM is trained (accuracy reported in Appendix) through the self-evolution process and then used to guide policy optimization. This configuration (e.g., green line) demonstrates consistent and significant performance improvement, approaching the results achieved with direct GT rewards.
The key finding from Figure 5 is that SEEA-R1, when utilizing its self-trained MGRM, can effectively learn and enhance its policy. This highlights the MGRM’s capability to provide reliable internal reward signals, enabling robust self-evolution and continuous performance gains without persistent dependence on external GT environment rewards.
Figure 5: Comparison of success rates of SEEA-R1 under different reward configurations.
4.2.4 Generalization on Out-of-Distribution Embodied Benchmark
To assess the generalization capabilities of proposed SEEA-R1, we conducted evaluations on the EmbodiedEval benchmark, an out-of-distribution (OOD) multi-modal embodied simulation test set. The success rates are reported in Table 3. Notably, our SEEA-R1 method, when fine-tuned with Tree-GRPO, demonstrates significantly improved performance after training. It achieves an overall success rate of 19.88%, which not only shows an uplift compared to the base model (Qwen2.5-VL-7B-Instruct at 18.29%) but also markedly outperforms models fine-tuned with SFT (16.77%) and DPO (16.77%). These results underscore the strong generalization performance of Tree-GRPO approach on this challenging, unseen benchmark.
Table 3: Success Rate in Embodiedeval benchmark
Model AttributeQA Navigation Object Interaction Social Interaction SpatialQA Overall
GPT-4o* achiam2023gpt 35.79 31.03 10.11 11.76 32.69 25.00
Qwen2.5-VL-7B-Instruct* 34.74 18.97 5.62 0.00 21.15 18.29
Qwen-VL-Max* bai2023versatile 37.89 24.14 24.91 8.82 17.31 21.04
LLaVA-NEXT-72B* zhang2024llava 23.16 12.07 7.83 0.00 5.77 10.67
LLaVA-OneVision-72B* zhang2024llava 26.32 10.34 1.12 0.00 19.23 12.80
Qwen2.5-VL-7B-Instruct + MCTS + SFT 28.42 20.69 4.49 5.88 19.23 16.77
Qwen2.5-VL-7B-Instruct + MCTS + DPO 27.37 17.24 5.62 5.88 23.08 16.77
SEEA-R1 30.85 21.05 4.60 6.25 32.69 19.88
5 Conclusion
We introduced SEEA-R1, the first framework to adapt reinforcement fine-tuning (RFT) for training self-evolving embodied agents. To address the challenges of sparse rewards and limited generalization for embodied domains, SEEA-R1 integrates Tree-GRPO, which leverages MCTS to densify reward signals, and MGRM, a multi-modal reward model that generalizes across tasks and environments. SEEA-R1 achieves new state-of-the-art performance on ALFWorld, outperforming previous models, including GPT-4o, and demonstrates strong generalization in self-learning settings. These results demonstrate the potential of SEEA-R1 as a scalable framework for embodied reasoning and autonomous learning.
Limitation Despite its strong performance and generalization, SEEA-R1, like current embodied AI, does not yet fully address the complexities of operating in highly dynamic and unpredictable real-world environments. This reflects a broader, significant challenge for the field, indicating a substantial path forward for future research to bridge the gap to truly autonomous real-world agents. Future work might explore scaling experiments involving more diverse environments, larger sample size, and bigger models, aiming to enhance SEEA-R1’s capacity for dynamic environments.
6 Acknowledgment
This work was supported by the National Natural Science Foundation of China (62476011).
Appendix A Appendix Overview
This appendix provides detailed supplementary material to complement our main paper. It is organized as follows:
  • Appendix B: Algorithm Details
    We outline the overall training procedure of our Self-Evolving Embodied Agent (SEEA-R1), as detailed in Algorithm 1.
  • Appendix C: Embodied Agent Formulation
    We describe the formalization of the embodied agent’s interaction with the environment as a Partially Observable Markov Decision Process (POMDP), detailing the observation, action, state, and reward components.
  • Appendix D: MCTS Implementation
    We elaborate on the Monte Carlo Tree Search (MCTS) mechanism, a core component of our SEEA-R1, covering its initialization, selection, expansion, simulation, and backup phases.
  • Appendix E: Exploration of World Model as Environmental Feedback
    We present our findings from exploring the use of a state-of-the-art world model (Kling) as environmental feedback in the self-evolving agent, discussing the challenges encountered such as visual hallucinations, limited instruction understanding, and generation inefficiency.
  • Appendix G: Dataset Details
    We describe the ALFWorld dataset used in our experiments, including its structure, evaluation splits, and the rationale behind focusing on the test-unseen setting to assess out-of-distribution generalization.
  • Appendix H: Extended Experimental Results
    This section presents comprehensive experimental results, including:
    • Evaluation on the General Multimodal Benchmark (MMBench) with a detailed breakdown of reasoning and perception capabilities.
    • Comparative performance across diverse multimodal benchmarks (MMStar, MMMU, HallusionBench, AI2D, and OCRBench).
    • An ablation study investigating the impact of sample size and batch size on training stability and success rate.
    • An algorithmic comparison demonstrating the superiority of our proposed Tree-GRPO over baselines like MCTS+DPO and MCTS+SFT.
    • Analysis of the long-term performance of Tree-GRPO over extended training iterations.
    • Investigation into the impact of iterative self-evolution on MCTS performance and generated data quality, including evaluations using Supervised Fine-Tuning (SFT).
  • Appendix I: Case Studies on ALFWorld Tasks
    We present representative execution trajectories across various ALFWorld task categories (e.g., pick-and-place, look, clean, temperature manipulation) to highlight the SEEA-R1 agent’s capabilities. Visualizations of policy and reward model interactions are also included.
  • Appendix J: Prompts Used
    We provide the full prompts used for both policy and reward models across different datasets (ALFWorld and EmbodiedEval), ensuring the reproducibility of our instruction-tuning pipeline.
Appendix B Algorithm
The overall training procedure of the Self-Evolving Embodied Agent is as follows (Algorithm 1).
1:  Initialize: Embodied Agent $\pi_{\theta_{old}}\leftarrow\pi_{\mathrm{ref}}$, MCTS parameters, Tree-GRPO hyperparameters
2:  for iteration $I=1,2,\dots$ do
3:     // Data Evolution: Experience Generation via MCTS
4:     Collect tree-structured experience data $\mathcal{D}_{\theta_{old}}=\emptyset$.
5:     for each episode do
6:        Generate MCTS-tree using the current agent $\pi_{\theta_{\mathrm{old}}}$.
7:        Extract triple set $(s_{t},a_{t,i},pr_{t,i})_{i=1}^{G}$ from MCTS-tree, and add to $\mathcal{D}_{\theta_{old}}$.
8:     end for
9:     // Model Evolution: Co-refining Policy Model and Reward Model
10:     Update reward model parameters by using the GRPO algorithm using $\mathcal{D}_{\theta_{old}}$:
11:     Update agent parameters $\theta$ by optimizing the Tree-GRPO objective $\mathcal{J}(\theta)$ using $\mathcal{D}_{\theta_{old}}$:
12:     $\theta\leftarrow\text{Tree-GRPO\_Update}(\theta_{\mathrm{old}},\mathcal{D}_{\theta_{old}},\mathcal{J},\pi_{\mathrm{ref}})$
13:     $\theta_{\mathrm{old}}\leftarrow\theta$
14:  end for
15:  Return: Optimized Embodied Agent $\pi_{\theta}$
Algorithm 1 Self-Evolving Framework Training Loop
Appendix C Embodied Agent Formulation
We model the long-horizon decision-making, multi-turn interaction embodied scenario as a Partially Observable Markov Decision Process (POMDP), formalized as a 7-tuple $(\mathcal{S},\mathcal{A},T,R,\mathcal{O},\Omega,\gamma)$. Here, $\mathcal{S}$ denotes the partially observable state space of the environment; $\mathcal{A}$ is the action space; $T(s_{t+1}|s_{t},a_{t})$ represents the conditional transition probability; $R(s_{t},a_{t})$ is the reward function or reward model; $\mathcal{O}$ is the observation space; $\Omega(o_{t+1}|s_{t},a_{t})$ is the conditional observation probability; and $\gamma$ is the discount factor. The main components of the Embodied Agent are described below:
Embodied Agent Observation $o_{t}\in\mathcal{O}$ is the observation received from the environment at timestep $t$ after the agent takes action $a_{t}$ in state $s_{t}$. In pure text-based scenarios, the observation is a natural language string, e.g., "Observation: The drawer 1 is closed.". In multimodal scenarios, the observation is an image representing the visual content of the environment from the current perspective.
[…]
Embodied Agent State $s_{t}\in\mathcal{S}$ is the partially observable environment state, derived from the agent’s historical exploration. It includes the entire history of interactions from the beginning of the task up to the current timestep, formalized as $s_{t}=\{s_{0},(a_{0},o_{0}),(a_{1},o_{1}),\dots,(a_{t-1},o_{t-1})\}$. The complete context $s_{t}$ carries clues about the environment state and preserves the reasoning trajectory, enabling the Embodied Agent to maintain task context memory and perform multi-turn reasoning and actions in partially observable scenarios. In ALFWorld, state $s_{t}$ includes initial state $s_{0}$ (system prompt describing environment rules and background; user prompt describing the current task goal and providing few-shot examples) and alternating actions $a_{t}$ and observations $o_{t}$ (agent-generated thoughts and actions, and environment-returned observations).
Embodied Agent Reward $r_{t}\in R$ is the reward received by the agent at timestep $t$ for taking action $a_{t}$ in state $s_{t}$. It can be provided by a reward model or a predefined reward function.
Appendix D MCTS
MCTS is a crucial component of SEEA-R1, comprising four core steps: selection, expansion, simulation, and backup.
Initialization: The initial state $s_{0}$ is set as the root node of the search tree.
Selection: Starting from the root node, the Embodied Agent traverses down the tree. At state $s_{t}$, it selects the child node with the maximum Upper Confidence Bound for Trees (UCT) value:
$a_{t}=\mathop{\arg\max}\limits_{a_{t,i}}\left[Q(s_{t},a_{t,i})+c\,\sqrt{\frac{\ln N(s_{t-1},a_{t-1})}{1+N(s_{t},a_{t,i})}}\,\right]$ (2)
where $c$ is the exploration constant, $Q(s_{t},a_{t,i})$ is the action value, $N(s_{t},a_{t,i})$ is the visit count for the action, and $a_{t,i}$ is the $i$-th available action at the current timestep $t$.
Expansion: After continuous selection reaches a leaf node $(s_{L},a_{L})$, the agent first executes this action to obtain an observation $o_{L}$ from the environment. Then, it creates a new non-leaf node $s_{L+1}=\{s_{L},(a_{L},o_{L})\}$. Finally, it simultaneously samples multiple natural language text segments to expand the new non-leaf node $s_{L+1}$, obtaining a set of available actions $\{a_{L+1,i}\}_{i=1}^{G}$.
Simulation: From the new leaf node $s_{L+1}$, the Embodied Agent performs multiple rollouts until termination—either the task is completed, the agent gives up, or the maximum search depth is reached. The complete trajectory of the $j$-th rollout is formalized as $s_{0}\rightarrow s_{1}\rightarrow\dots\rightarrow s_{L+1}\rightarrow s_{L+2}^{(j)}\rightarrow\dots\rightarrow s_{T}^{(j)}$.
Backup: The results of multiple rollouts are used to update the visit counts $N(s_{t},a_{t})$, returns $R^{(j)}(s_{t},a_{t})$, and action values $Q(s_{t},a_{t})$ along the trajectory, starting from the terminal node and moving upwards. The update formulas are as follows:
$N(s_{t},a_{t})\leftarrow N(s_{t},a_{t})+1$ (3)
$R^{(j)}(s_{t},a_{t})=r(s_{t+1}^{(j)})+\gamma\,r(s_{t+2}^{(j)})+\dots+\gamma^{T-t-1}\,r(s_{T}^{(j)})$ (4)
$Q(s_{t},a_{t})=\mathbb{E}_{\begin{subarray}{c}a_{t}\sim\pi_{\theta_{\mathrm{old}}}(\mathcal{A}|s_{t})\end{subarray}}\bigl{[}R(s_{t},a_{t})\bigr{]}=\frac{\sum_{j=1}^{N(s_{t},a_{t})}R^{(j)}(s_{t},a_{t})}{N(s_{t},a_{t})}$ (5)
Figure 6: Pipeline of the experiment for the World Model as environmental feedback in Self-evolving Agent
Appendix E Exploration of World Model as environmental feedback in Self-evolving Agent
In a self-evolving agent, environmental feedback is critical and can take many forms—such as a simulated environment, a world model, or the real world. Recent research increasingly trains world models on large-scale data so that they can emulate reality. To test whether updated world models are sufficiently capable of serving as environmental feedback, we intend to train a self-evolving agent not only in a simulator but also within a world-model environment. For short, we use kling, a sota world model, as a world simulator in our experiment setting.
During this data-collection phase we found that when the agent relies on a world model for feedback, obtaining a single accurate and complete trajectory is extremely difficult. This suggests that current world models are still far from being fully reliable world simulators and have a long way to go.
We have identified the following aspects that require improvement in world models:
1. Frequent visual hallucinations in predicted videos. Hallucinations in the video-based world model are clichéd problems, and there have been many previous studies on how to solve them. In our data collection phase, even when we supply detailed prompts—and likewise with simpler prompts—to the sota world model, we often encounter hallucinations. Typical failures include physically impossible dynamics, objects that suddenly warp or vanish, or the spontaneous appearance of extra robot arms and other artifacts. Repeated tests confirm that the current generation of world models cannot reliably prevent such hallucinations; therefore, the resulting videos frequently fail to meet our requirements and make the model unsuitable as a dependable world simulator.
2. Limited understanding of task instructions. Misinterpreting an instruction is distinct from hallucination: the video looks coherent, yet it fails to follow the given command. We observe two main failure modes. (i) Generalisation gaps: when the model encounters an unfamiliar instruction or action, it often produces a static scene or executes the wrong motion—without hallucinating, but still violating the instruction. (ii) Planning errors: if the upstream planner issues a physically unsound command (e.g., “place the plate on the very edge of the table”), a faithful simulator should show the plate toppling. Today’s world models cannot reproduce such plausible but incorrect plans, revealing a deeper weakness in instruction comprehension.
3. Inefficient video generation. Large-scale visual diffusion models typically require 10–30 minutes to produce a 5–10-second clip. This lengthy turnaround severely slows data collection; if the output is unusable, we must restart the process and wait again. Consequently, training a self-learning agent with a world model demands substantially more time and compute than training in a conventional simulator.
Appendix F Code, Dataset and Extra Experiments
We include the detailed dataset construction and formulation in the supplement materials, together with more fine-grained ablation experiments on the model and data.
Figure 7: Performance comparison of SEEA-R1 across different tasks in ALFWorld over training iterations. Left: Success rate across tasks. Right: Average number of steps taken to complete tasks.
Appendix G Dataset
ALFWorld: The ALFWorld dataset is structured into a training set comprising 3321 games and a test set, further partitioned into test-seen (140 games) and test-unseen (134 games) splits. This distinction is crucial for assessing out-of-distribution (OOD) generalization, as unseen tasks introduce novel environments and instructions. For our experiments, we specifically evaluate our model on the test-unseen split. This focus on unseen scenarios is motivated by the need to rigorously assess the agent’s capacity for effective long-horizon planning and generalization to novel embodied tasks, which is critical for real-world deployment. ALFWorld’s embodied setup, sparse rewards, and requirement for OOD generalization collectively make it a rigorous benchmark for evaluating agents intended for complex real-world task-solving.
Appendix H Experiments
H.1 Evaluation on the Genenal Multi-Modal Benchmark
MMBench is a multimodal benchmark that subdivides reasoning and perception capabilities into six Level-2 dimensions: Logic Reasoning (LR), Attribute Reasoning (AR), Relation Reasoning (RR) for Reasoning, and Fine-Grained Perception-Single Instance (FP-S), Fine-Grained Perception-Cross Instance (FP-C), and Coarse Perception (CP) for Perception. The overall average score is derived from the average of scores for all 20 Level-3 capabilities, and Level-2 capability dimension scores are the average of all Level-3 capability scores within that dimension. Ground truth labels for the test set are not publicly available. The evaluation results are reported in Table 4.
Table 4: Performance comparison of various Multimodal Large Language Models on the MMBench benchmark.
Model Overall LR AR RR FP-S FP-C CP
Gemini-2.5-Pro 88.5 89.1 87.8 83.7 94.2 94 83.7
GPT-4o 86.8 87.5 91.4 83.3 84.9 91 85.4
InternVL3-8B 82.3 71.2 84.1 82.9 88.2 81.3 80.9
Qwen2.5-VL-7B 82.4 71.7 84.9 80.2 89.8 80.1 81.3
LLaVA-OneVision-7B 80.9 67.9 88.2 82.2 84.9 73.8 81.9
SEEA-R1 81.4 66.9 82.9 81.0 90.3 75.4 81.7
To comprehensively evaluate the capabilities of Multimodal Large Language Models (MM-LLMs), we report performance across a suite of diverse benchmarks in Table 5, each targeting distinct aspects of multimodal intelligence. MMStar assesses general multimodal reasoning and understanding. MMMU (Massive Multi-discipline Multimodal Understanding) focuses on complex, expert-level multimodal reasoning across various academic disciplines. HallusionBench specifically targets the detection and mitigation of hallucinations in MM-LLMs, particularly when dealing with potentially misleading visual information. AI2D (AI2 Diagrams) evaluates the models’ ability to understand and reason about scientific diagrams, involving both visual perception and logical inference. Lastly, OCRBench measures the models’ optical character recognition capabilities, evaluating their effectiveness in extracting and understanding text embedded within images. This comprehensive evaluation suite provides a holistic view of each model’s strengths and limitations in real-world multimodal applications.
Table 5: Performance comparison of various Multimodal Large Language Models on diverse multi-modal benchmarks.
Model MMStar MMMU HallusionBench AI2D OCRBench
Gemini-2.5-Pro 73.6 74.7 64.1 89.5 862
GPT-4o 70.2 72.9 57 86.3 822
SenseNova-V6-Pro 73.7 70.4 67.1 89.2 895
InternVL3-8B 68.7 62.2 49 85.1 884
Qwen2.5-VL-7B 64.1 58 51.9 84.3 888
LLaVA-OneVision-7B 56.7 46.8 47.5 82.8 697
SEEA-R1 60.2 48.5 66.4 79.4 766
Figure 8: Performance comparison across different sample and batch sizes.
H.2 Impact of Sample Size and Batch Size on Training Stability
We investigate how training configurations affect model performance. As illustrated in Figure 8, larger sample and batch sizes significantly influence the final success rate on the AlfWorld unseen test set. Specifically, the largest configuration (Sample=1024, Batch=512) consistently demonstrates a clear performance advantage, achieving a final success rate of 41.79%. In stark contrast, while the medium setting (Sample=512, Batch=128) reached a peak success rate of 32.46% during its intermediate iterations, and the smallest configuration (Sample=128, Batch=32) similarly achieved its highest accuracy of 25.75% at an earlier iteration, their overall final performance was notably lower and exhibited less stable convergence compared to the largest configuration. These results collectively indicate that increasing the sample and batch sizes contributes to more stable and effective policy updates in GRPO-based training, likely due to reduced variance in gradient estimation, ultimately leading to superior and more consistent performance on unseen tasks.
Figure 9: Performance comparison of Tree-GRPO against MCTS+DPO and MCTS+SFT across training iterations.
H.3 Comparison with Different Algorithms
We investigate the comparative performance of Tree-GRPO against two established baseline methods: MCTS integrated with DPO (MCTS+DPO) and MCTS with SFT (MCTS+SFT). As illustrated in Figure 9, Tree-GRPO consistently outperforms both baselines across training iterations on the AlfWorld unseen test set. Tree-GRPO reached a peak success rate of 32.46%, which is significantly higher than the 25.37% peak achieved by MCTS+DPO and the 16.79% peak of MCTS+SFT.
This superior performance can be largely attributed to GRPO’s on-policy nature, which facilitates more accurate and stable policy improvement through direct, interaction-aligned updates. In contrast, DPO is fundamentally an off-policy algorithm; it shares practical limitations with SFT, as both methods heavily rely on static, pre-collected data and lack the dynamic optimization benefits derived from online rollouts. Consequently, the off-policy characteristic of DPO contributes to its observed performance ceiling, while the stagnation and eventual degradation in the performance of MCTS+SFT clearly reflect the inherent limitations of purely supervised learning in complex interactive environments. The on-policy nature of GRPO, especially when combined with tree-based exploration, enables more stable and effective updates, leading to superior sample efficiency and enhanced robustness in reinforcement learning.
Figure 10: Learning curve of Tree-GRPO over 30 training iterations
H.4 Long-Term Performance of Tree-GRPO
To thoroughly assess the long-term effectiveness of Tree-GRPO, we extended the training duration of the model to 30 iterations. As illustrated in Figure 10, the success rate on the AlfWorld unseen test set demonstrates a consistent and substantial increase, rising from an initial 11.57% to a robust final success rate of 39.48%. While minor fluctuations are observed, which are indicative of exploration-induced variance inherent in reinforcement learning, the overall upward momentum remains undeniably consistent. This sustained improvement over an extended training horizon unequivocally confirms the scalability and strong convergence potential of Tree-GRPO. Such reliable long-term performance makes Tree-GRPO particularly well-suited for practical deployment in complex, long-horizon decision-making tasks where consistent improvement and robustness are paramount.
H.5 Impact of Iterative Self-Evolution on MCTS Performance and Data Quality
This experiment investigates how iterative self-evolution affects Monte Carlo Tree Search (MCTS) performance and the quality of trajectories it generates. We hypothesize that refining the MCTS’s underlying model through successive self-evolution iterations leads to progressively higher-quality sampled trajectories, as measured by increased average rewards.
Figure 11: Iterative Self-Evolution Enhances MCTS Performance and Sampled Trajectory Quality. Each line represents a distinct self-evolution iteration (from iter_1 to iter_12), plotting the average cumulative maximum goal reward against MCTS search steps. The color gradient (blues for early iterations, reds for later ones) highlights the progression.
The results are presented in Figure 11. A clear positive trend is evident: later self-evolution iterations (warmer colors, e.g., iter_12) consistently achieve higher average cumulative maximum rewards compared to earlier iterations (cooler colors, e.g., iter_1) across various MCTS search step counts. For instance, at 15 MCTS search steps, iter_12 attains an average reward of approximately 0.8, whereas iter_1 reaches about 0.7. This performance gap tends to widen with an increasing number of search steps, implying that models from later iterations benefit more substantially from deeper MCTS searches.
This observed upward trend in average reward across self-evolution iterations strongly suggests that the models become more proficient at guiding MCTS towards successful outcomes. Consequently, trajectories sampled by MCTS using these more advanced models are of increasingly higher quality. This improvement in sampled data is crucial, as it provides a more effective training signal for subsequent model refinement, thereby validating the efficacy of the self-evolutionary approach in enhancing both agent performance and data generation.
H.6 Evaluating Data Quality from Iterative MCTS + GRPO via Supervised Fine-Tuning
To assess the evolution of data quality throughout the Iterative MCTS + GRPO training process, we conducted an auxiliary experiment. We aimed to determine if data collected in later iterations progressively improves and if this improvement translates to enhanced performance when a separate, pre-trained model (Qwen2.5-VL-7B-Instruct) is fine-tuned using this data.
For this experiment, data generated by the policy model after each iteration of Iterative MCTS + GRPO was collected. These samples were then filtered, retaining only those with an Advantage value greater than zero, which we considered as potentially high-quality instances. Subsequently, this filtered data from each respective iteration was used for Supervised Fine-Tuning (SFT) of the Qwen2.5-VL-7B-Instruct model. We hypothesized that if data quality improves with more iterations, the SFT performance of Qwen2.5-VL-7B-Instruct would exhibit a corresponding upward trend when trained on data from later iterations.
Figure 12: Success Rate Comparison: Iterative MCTS + GRPO vs. SFT on Generated Data. The orange line tracks the success rate of the Iterative MCTS + GRPO policy over its training iterations. The blue bars represent the success rate of the Qwen2.5-VL-7B-Instruct model after SFT using data collected from the corresponding Iterative MCTS + GRPO iteration (filtered for Advantage > 0).
However, as depicted by the blue bars in Figure 12, the SFT performance of Qwen2.5-VL-7B-Instruct did not show a monotonically increasing trend with the iteration number of the source data. Instead, the success rate fluctuated. For instance, SFT on data from iteration 3 yielded a success rate of approximately 0.247, but this dipped with data from iteration 4 (approx. 0.164) and iteration 5 (approx. 0.179). While performance peaked again with data from iteration 7 (approx. 0.268), it subsequently declined for data from later iterations (8, 9, and 10).
This observation—that SFT performance on a fixed baseline model does not consistently improve with the iteration number of the data-generating policy—suggests a nuanced aspect of iterative reinforcement learning. While the Iterative MCTS + GRPO process itself (orange line in Figure 12) demonstrates clear improvement in its own success rate, the data generated at each step may become increasingly specialized or "tailored" to the current state and capabilities of the policy model that produced it. Consequently, this data, while optimal for the self-improvement of the generating policy, may not be universally or increasingly beneficial for fine-tuning a different, general-purpose model like Qwen2.5-VL-7B-Instruct. This suggests that the data collected by the evolving policy becomes more "self-serving" and less generalizable for enhancing distinct models as training progresses
Appendix I Case Studies on AlfWorld Tasks
Figure 13: AlfWorld Task Demonstrations (a–d). (a) Put a pencil in shelf — step-by-step object manipulation. (b) Put two CDs in safe — compound goal handling. (c) Look at object — spatial and property-based reasoning. (d) Heat mug and place in coffeemachine — multi-step planning.
Figure 14: AlfWorld Task Demonstrations (e–h). (e) Put hot tomato in trash — handling state change before disposal. (f) Cool an object — inverse state manipulation. (g) Clean the egg and microwave — sequential object preparation. (h) Pick and Look — combined manipulation and inspection.
We present representative execution trajectories across various task categories in the AlfWorld benchmark. Each case highlights the SEEA-R1 agent’s ability to understand language instructions, plan goal-directed behaviors, and interact with the environment effectively.
I.1 AlfWorld: Pick and Place Task
Figure 13 illustrates a typical Pick and Place task, where the agent is instructed to place a pencil into a shelf. The agent must first locate and pick up the correct object, navigate to the target receptacle, and complete the placement action. This requires not only spatial reasoning but also precise object manipulation.
I.2 AlfWorld: Multi-Object Placement Task
A more complex variant of the pick and place paradigm is shown in Figure 13, where the agent is asked to place two CDs into a safe. This scenario introduces increased goal complexity, requiring multi-object tracking, inventory management, and the satisfaction of compound task conditions.
I.3 AlfWorld: Look Task
As shown in Figure 13, the Look task evaluates the agent’s perceptual capabilities. The instruction "Look at CD under the desklamp" requires SEEA-R1 to locate the correct spatial reference (under the desklamp), identify the target object, and perform an observation action. This task emphasizes the importance of grounded language understanding and fine-grained spatial perception.
I.4 AlfWorld: Temperature Manipulation Tasks
Figures 13 and 14 illustrate representative instances of Heat tasks in the AlfWorld benchmark. In these scenarios, the agent is required to manipulate the thermal state of target objects—such as heating a mug prior to placing it in a coffee machine, or ensuring that a tomato is sufficiently warmed before disposal. Successfully completing these tasks demands the agent to not only identify the correct object and comprehend the temporal order of subgoals (e.g., "heat before use"), but also to correctly interact with relevant environmental appliances (e.g., microwaves or stoves). These tasks pose unique challenges in reasoning about object affordances, causal relationships, and multi-step transformations, as the agent must infer the implicit preconditions necessary for goal satisfaction and execute a valid sequence of actions accordingly.
I.5 AlfWorld: Clean Task
In Figure 14, the Clean task involves transforming the object’s state (e.g., cleaning an egg before microwaving it). This category tests the agent’s ability to recognize when preprocessing steps are required to satisfy high-level goals.
I.6 AlfWorld: Composite Pick & Look Task
Figure 14 presents a composite task that blends object manipulation with perception. The agent must pick up an item and inspect its surroundings, requiring coordination of multiple sub-skills within a single instruction. This reflects real-world complexity and cross-modal reasoning demands.
I.7 Visualization of Policy and Reward Model Interactions
Figure 15 provides qualitative examples of the policy and reward models interacting within the environment. Each panel shows the visual state, the reward model’s reasoning (‘RM: <think>…‘), and the policy model’s chosen action (‘Actor: Action:…‘), illustrating their step-by-step interplay.
Figure 15: Policy and reward model interactions. Shows state, reward model thoughts, and actor actions step-by-step.
Appendix J Prompts
J.1 Prompt for the ALFWorld Dataset
[…]
J.2 Prompt for the EmbodiedEval Dataset
[…]
J.3 Reward Model
Instruction Prompt for Reward Model A conversation between User and Assistant. The user provides the task and the current state, and the Assistant evaluates the state and determine if the current state is success, failure, or continue. Failure means the user cannot achieve success through further actions, while continue means the user can still achieve success through additional actions.  The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think> <answer> Success/Failure/Continue </answer>.